Opened 11 years ago

Last modified 6 years ago

#362 closed flight processing

EUFAR10/07, flight day 234d, Summer School Geo Site — at Version 20

Reported by: iopa Owned by: anch
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: ~arsf/workspace/EUFAR10_07-2010_234d_Summer_School_Geo_Site

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of lithology, fault strucuture,
erosion parameters and soil components

Priority: a4h

PI: A. Zlinszky

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

Sensors:

  • Eagle (04/03/2011)
  • Hawk (04/03/2011)
  • Leica LIDAR (Requested by PI)
  • Digital Photography (04/03/2011)

Change History (20)

comment:1 Changed 11 years ago by iopa

  • Description modified (diff)

comment:2 Changed 11 years ago by knpa

  • Description modified (diff)

Added scientific objective.

comment:3 Changed 11 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:4 Changed 11 years ago by benj

  • Description modified (diff)

comment:5 Changed 11 years ago by benj

  • Description modified (diff)

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

comment:6 Changed 11 years ago by emca

Processed the camera imagery from raw to tiff format

comment:7 Changed 11 years ago by anch

Applanix has been processed for 238e which covers this flight, so proc has been copied into the applanix folder from 238e

comment:8 Changed 11 years ago by anch

Starting hyperspectral processing

comment:9 Changed 11 years ago by anch

Submitted hyperspectral for delivery check

comment:10 Changed 11 years ago by anch

Starting LiDAR processing.

comment:11 Changed 11 years ago by anch

IPAS from 234e covers this flight so I have copied it from there.

comment:12 Changed 11 years ago by iopa

Checking hyperspectral.

comment:13 Changed 11 years ago by iopa

Hyperspectal checking done;

  • there's been a keyboard slip somewhere, proj code is misspelt in delivery directory, DEM name, logsheet, and DEM name within the readme.
  • eagle/hawk flightline id's in readme for line 3 should be 012

else the data seems good and ready to deliver.

comment:14 Changed 11 years ago by emca

Starting camera processing

comment:15 Changed 11 years ago by anch

Submitted lidar for delivery check

comment:16 Changed 11 years ago by emca

Camera ready for delivery check at
~airborne/workspace/EUFAR10_07-2010_234d_Summer_School_Geo_Site/delivery/20110303

comment:17 Changed 11 years ago by anch

Checking camera

comment:18 Changed 11 years ago by anch

Camera delivery is fine and ready to deliver

comment:19 Changed 11 years ago by knpa

  • Owner set to anch
  • Status changed from new to assigned

Taking lidar delivery check.

comment:20 Changed 11 years ago by anch

  • Description modified (diff)

Hyperspectral, and camera sent to PI

Note: See TracTickets for help on using tickets.